News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 30 June 2000 3.30pm EDT

During the last week, Chandra completed the observing schedule as planned.

The HETG was inserted via hardware commands at 8.30pm EDT on Jun 25. The insertion was nominal as expected, with the insert limit switch functioning to cut power to the MCE. The HETG was retracted using timed hardware commanding at 10.30pm on Jun 26. The retraction was timed for 155 seconds and telemetry indicated a nominal move. This duration leaves the HETG just off the bumper spring in the fully retracted position. A technical meeting was held on Jun 27-28 at the OCC to discuss OTG data, the fault tree, and options for restoring full grating operations via a flight software patch. Results from the meeting were encouraging and a plan for resuming short term HETG operation from the daily load (without patching flight software), and limited LETG operation via real-time commanding was discussed. A flight software patch to return full operation is under assessment.

A real-time procedure was run at 9pm EDT on June 28 to power cycle the ACIS Front End Processors (FEPs) after science data processed from June 26 indicated anomalous data from the chips connected to FEP-0 and FEP-3. (The prime science target is always read by FEP-1.) The procedure returned the FEP's to normal operation. The data are consistent with "Threshold-Plane Lock-Up" of two of the FEPs which can occur when the FEP attempts to perform multiple functions at once (in this case 3). A very unusual combination of ACIS parameter settings had been used in the squeegee mode test proceeding the lock-up and is almost certainly the cause. These settings are not planned for use in the future and ground tests are in work to confirm the cause.

The planned observations for the coming week are as follows.

------------------------------------------
E0102-72.3 (multiple)  ACIS-S          Jul  3
E0102-72.3 (multiple)  ACIS-I          Jul  3
Radiation Belt                         Jul  4
Ophiuchi Mosaic #10    ACIS-I
Ophiuchi Mosaic #8     ACIS-I
NGC 6440               ACIS-S
VB 10                  ACIS-S
G21.5-0.9 (multiple)   ACIS-S          Jul  5
G21.5-0.9 (multiple)   ACIS-I          Jul  6
Radiation Belt
Abell 1423             ACIS-I
NGC 4314               ACIS-S          Jul  7
CB 58                  ACIS-S
HD 135379              ACIS-S
1736-297 Van der       ACIS-S
Galactic Center Arc    ACIS-I          Jul  8
W49B                   ACIS-S
Radiation Belt                         Jul  9
HCG 92                 ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports